OpenStudy (anonymous):

Help me on how to solve these questions with complex numbers: a) 2-i / sqrt2 - i ; b) 3 + 4i / 4 - 3i ; c) 3 / 1 + i ; d) 4i / 2 - 3i

OpenStudy (phi):

if you have A/B where B is complex, multiply top and bottom by the complex conjugate of B \[ \frac{A}{B} \cdot \frac{B^*}{B^*} \]

OpenStudy (anonymous):

I only have one question. With the square root though, how do I simplify that?

OpenStudy (phi):

you don't. for (a) you have \[ \frac{2-i}{\sqrt{2}-i}\cdot \frac{\sqrt{2}+i}{\sqrt{2}+i} \\ \frac{(2-i)( \sqrt{2}+i)}{2-i^2}\\ \frac{(2\sqrt{2}+1) +(2-\sqrt{2})i}{3} \]

OpenStudy (anonymous):

okay thanks. Also, I got the answer -10/25 + i for b) but I don't think it's the correct solution.

OpenStudy (phi):

how ?

OpenStudy (anonymous):

I have my insecurities, haha I just want to make sure I understand complex numbers since I have a test next week.

OpenStudy (phi):

how did you get your answer for (b)? It looks wrong.

OpenStudy (anonymous):

The complex conjugate. I must have done something wrong in the second last or last step.

OpenStudy (phi):

the bottom is 25 what do you get for the top.... can you show the details as you multiply (3 + 4i ) * (4 + 3i )

OpenStudy (anonymous):

 first step